Obtain the corresponding potential energy function of the intermolecular force F(r) = -ar^-7 + br^-9


1
Expert's answer
2021-06-07T09:31:57-0400

Answer

Potential energy can be given as

"U=-F(r) dr\\U=-(-ar^{-7}+br^{-9}) dr\\U=\frac{-ar^{-6}}{6}+\frac{br^{-8}}{8}"